OpenStudy (anonymous):

The sum of two numbers is 70. One number is 4 times the other. Find the numbers. @geerky42

geerky42 (geerky42):

we need to create two variables. a and b? sure. So we have a + b = 70 and a = 4b

geerky42 (geerky42):

we need to find a and b. use substitution method

geerky42 (geerky42):

can you do it?

OpenStudy (anonymous):

4b+b=70

geerky42 (geerky42):

yep so b = ??

OpenStudy (anonymous):

5b=70 b=14!

OpenStudy (anonymous):

4b+14=70

geerky42 (geerky42):

correct. so we know one of them are 14. what's the other?

geerky42 (geerky42):

we found value for b. now we need to know a

OpenStudy (anonymous):

a+14=70 a=56

geerky42 (geerky42):

excellence! so that's your answer. 14 and 56

OpenStudy (anonymous):

Thank you so much! Can you help me with the others too? I have to pass this quiz :/

geerky42 (geerky42):

sure. create new question. I am going to leave soon. sorry...